# Joint Multimodal Transformer for Emotion Recognition in the Wild

### v2

@@ Abstract
Systems for multimodal emotion recognition (MMER) can typically outperform unimodal systems by leveraging the inter- and intra-modal relationships between, e.g., visual, textual, physiological, and auditory modalities. In this paper, an MMER method is proposed that relies on a joint multimodal transformer for fusion with key-based cross-attention. This framework aims to exploit the diverse and complementary nature of different modalities to improve predictive accuracy. Separate backbones capture intra-modal spatiotemporal dependencies within each modality over video sequences. Subsequently, a joint multimodal transformer fusion architecture integrates the individual modality embeddings, allowing the model to capture inter-modal and intra-modal relationships effectively. Extensive experiments [^1] Code: https://github.com/PoloWlg/Joint-Multimodal-Transformer-6th-ABAW. on two challenging expression recognition tasks: (1) dimensional emotion recognition on the Affwild2 dataset (with face and voice), and (2) pain estimation on the Biovid dataset (with face and biosensors), indicate that the proposed method can work effectively with different modalities. Empirical results show that MMER systems with our proposed fusion method allow us to outperform relevant baseline and state-of-the-art methods.
@@
Figure 1: (Top) Illustration of the vanilla multimodal transformer fusion architecture. (Bottom) Multimodal transformer fusion with joint representation (in red) for two input sources, A and B.
@@ 1 Introduction
Human-computer interaction is applied in a wide range of real-world scenarios, e.g., health care, IoT, and autonomous driving. Researchers have classified human emotions in different ways, most notably according to discrete categories, ordinal intensity levels, and the valence/arousal circumplex [1]. With the recent advancements in deep learning and sensor technologies, research in affective computing has evolved from lab-controlled to real-world (in the wild) scenarios. In the latter, human emotions are usually expressed over a broader spectrum beyond the six basic categorical expressions - anger, disgust, fear, happy, sad, and surprise [9]. Therefore, there is much interest in analyzing and modeling complex and subtle expressions of emotions in real-world scenarios. For instance, the spectrum of emotions can be formulated as dimensional emotional recognition (ER), where complex human emotions are represented along the axes of arousal (intensity) and valence (positiveness).
@@
Several methods have been proposed for video-based ER and recurrent networks have been employed to capture the intra-modal temporal dependencies from video sequences [50, 28, 5]. Recently, attention-based methods have been introduced to extract features that are the most relevant to downstream tasks. Cross-attention-based methods have also been [48] employed to capture the inter-modal association between the audio, visual, and other modalities. Lu et al. [35] proposed ViLBERT, which was the seminal work in multimodal co-attention. Since then, many transformer-based cross-attention methods have been proposed [43, 57]. These methods, however, cannot effectively capture the intra-modal temporal dynamics. Further, they specialize in capturing the complementary information among the modalities but do not have the mechanism to capture the redundant information explicitly.
The proposed method introduces the third branch with the joint representation of the multiple modalities, as shown in Figure 1. By incorporating a joint representation branch, the model can access additional contextual information that might not be fully captured by cross-attention alone. This can help improve the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences. This helps in mitigating the sensitivity of cross-attention to noisy inputs and improves the overall performance of the system.
@@
(1) This paper proposes a joint multimodal transformer fusion framework that leverages joint modality representations. It captures inter- and intra-modal information in videos using key-based cross-attention. (2) A joint multimodal transformer multimodal fusion that exploits the redundant and complementary associations among modalities. (3) An extensive set of experiments on the challenging emotion recognition (pain estimation on BioVid and dimension valence-arousal assessment on Affwild2) datasets indicate that our proposed multimodal fusion architecture can outperform relevant baseline and state-of-the-art methods.
@@ 2.1 Multimodal Methods
MMER refers to the integration of multiple sources of information (modalities) to improve the accuracy and robustness of automated emotion recognition systems, at the expense of complexity. These modalities typically include visual, audio, textual, and physiological. The seminal work in multimodal deep learning was proposed by Ngiam et al. [39], where the features from the audio and visual modalities were extracted separately, and then autoencoders and Restricted Boltzmann Machines were used to feature fusion. Tzirakis et al. [55] proposed one of the early approaches for A-V fusion for dimensional emotion recognition, in which the visual features were extracted using a ResNet50 and the audio features were obtained using a 1D convolutional neural network (CNN). The modality-specific features were concatenated and fed to a recurrent net for simultaneous temporal modeling and modality fusion. An empirical study was presented by Juan et al. [42], where the authors studied the impact of fine-tuning multiple layers in a pretrained CNN for the visual modality. A two-stream autoencoder with a long short-term memory (LSTM) network was proposed by Nguyen et al. [40] to jointly learn and compact representative features from the visual and audio modalities. A knowledge distillation-based approach was investigated by Schonevald et al. [50] for visual modality. For the audio modality, spectrograms were obtained and fed to a CNN model, and the two modalities were fused using a recurrent net. A novel self-distillation scheme was put forward by Deng et al. [6] to overcome the problem of noisy labels in a multitasking setting. A two-stream aural visual (TSAV) network was proposed by Kuhnke et al. [28], in which the audio features were extracted using a ResNet18, and the visual features were extracted using a 3D-CNN. The obtained embeddings were fed to a specially designed TSAV network for information fusion.
Pain classification is one of the primary problems in affective computing. Researchers have proposed many multimodal datasets for the pain estimation task. The facial activity descriptors method for pain estimation was introduced by Werner et al. [59]. Dragomir et al. [7] propose a subject-independent method from facial images with a residual learning technique. A Sparse LSTM-based method was proposed by Zhi et al. [65] to solve the problem of vanishing gradients in temporal learning. Morabit et al. [38] proposed a data-efficient image transformer. To process multiscale electrodermal activity signals, a SE-Net-based network was proposed by Lu et al. [36]. Multimodal solutions to fuse the physiological and visual modalities were proposed by Werner et al. [59], Kachele et al. [13], and Zhu et al. [65]. Physiological signals are more discriminative for pain classification than the visual modality.
@@ 2.2 Attention-Based and Transformer Methods
Since its inception, attention models have shown extraordinary performance in many applications. These models have been extensively investigated for capturing the inter and intra-modal associations between the audio and visual modalities for tasks like action localization [32], A-V event localization [8], and multimodal emotion recognition [43]. An attention-based fusion mechanism was proposed by Zhang et al. [64], 3D-CNNs and 2D-CNNs were used to extract multi-features in the visual modality, and for the audio modality, a 2D-CNN was used to learn representation from spectrograms. Specialized scoring functions were used to re-weight the audio and visual features. Recently, cross-modal attention has shown promising results because of its ability to model inter-modal relationships. Srinivas et al. [43] explored a transformer network with encoder layers, where cross-modal attention is used to fuse audio and visual features for continuous arousal/valence prediction in the wild. Tzirakis et al. [54] explored the idea of cross-attention in conjunction with self-attention. The authors proposed a transformer-based fusion architecture. Although the methods mentioned above have used cross-modal attention with transformers, they do not have any explicit mechanism to capture semantic relevance between the A-V features, particularly the intra-modal correlations. Zhang et al. [63] proposed a method for A-V fusion using leader-follower attentive fusion for continuous arousal/valence prediction. Attention weights are combined with the encoded visual and audio features. Cross attention presented in Praveen et al. [48] has shown a substantial increase in performance by using cross-correlation across the individual features. In contrast, our proposed method uses key-based cross-attention in multimodal transformers and explores the idea of feeding the joint A-V feature vector. By feeding the joint A-V feature representation, the proposed method effectively captures the inter- and intra-modal relationships simultaneously by interacting across itself and the other modalities.
@@
Huang et al. [11] investigated the idea of multi-head attention in transformer-based fusion architecture, which was further combined with LSTM to capture the high-level representations. Tran et al. [53] proposed a cross-modal transformer architecture that consisted of a multimodal cross-modal attention block, where the Queries were generated from one modality and the key-values were generated from the other modality. Le et al. [31] put forward an end-to-end transformer-based fusion mechanism for multilabel multimodal emotion classification; the model consisted of three parts: i) three backbone networks for visual, audio, and textual feature extractor, ii) a transformer network for information fusion, and iii) classification network. Zhou et al. [66] proposed a transformer-based fusion scheme along with the temporal convolutional network (TCN); the audio and visual features were extracted using pretrained backbones followed by a TCN, the output of TCN was concatenated and fed to a transformer encoder block. A multilayer perceptron (MLP) was then used for the final prediction.
@@
All the aforementioned transformer-based fusion architectures primarily focus on intermodality correlation. In contrast, the proposed method, in addition to modeling the intermodality relationships to capture the complementarity between modalities, we explicitly feed the joint (combined) features to the multimodal transformer to introduce redundancy. By incorporating this third joint representation branch, the proposed model has access to enhanced contextual information that might not be fully captured by cross-attention alone. Doing this improves the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences. Having this third joint representation allows the model to dynamically focus on this newly introduced information in sequences where both modalities are noisy simultaneously. This helps in mitigating the sensitivity of cross-attention to noisy inputs and improves the overall performance of the system.
@@ 3 Proposed Approach
The proposed method is a hierarchical fusion mechanism, where the intra-modality features are combined using transformer-based self-attention, and cross-modality features are fused using transformer-based cross-attention. Further, we feed a third joint representation to the joint transformer module (JTM) to enhance robustness. The $K$ (key matrix), $V$ (value matrix), and $Q$ (query matrix) vectors are shared among the six transformer blocks. In the end, the output of these six blocks is again fed to a transformer self-attention block to dynamically weigh the most relevant representations. The final prediction is made using fully connected (FC) layers.
@@ 3.1 Modality Specific Feature Extraction
In the first step, modality-specific features are extracted using backbones. The proposed method allows to combine multiple backbones for each modality to improve system robustness. The extracted feature vectors from each backbone are fed to a transformer self-attention block. The combined feature vector is the representation of the particular modality. For example, to capture information about a person’s emotional state, we can use an R(2+1)D CNN pretrained on the Kinetics-400 dataset [29] to extract visual features. For the audio modality, we could extract features using a ResNet18 [10] CNN with a GRU [4].
@@ 3.2 Multi-transformer Attention Fusion
Each representation is then fed to a specific encoder. Our model is composed of three different encoders, one for each type of feature ${\bm{F}}_{\mathbf{B}}$, ${\bm{F}}_{\mathbf{A}}$, and $\bm{J_{AB}}$. Each encoder consists of a multi-head-self-attention (Eq. 2) followed by a fully connected feed-forward network. Residual connection and layer normalization are performed around both of these layers. The key is used to associate a sequence to a key value, the value matrix holds information that is ultimately used to compute the output of the attention mechanism, and the query matrix represents a set of vectors used to query the key-value pairs. The ${K}$, ${V}$, ${Q}$ matrix are calculated this way: ${K={X}{W_{K}}}$, ${Q={X}{W_{Q}}}$, ${V={X}{W_{V}}}$. ${X}$ corresponds to one of the sources of information and ${W_{K}}$, ${W_{Q}}$, ${W_{V}}$ are the weights of the key, query, value matrices respectively. The output values of the self-attention layers are given by:
${Attention(Q,K,V)=softmax({K}{Q^{T}}/{d_{k}})V}$ (2)
@@
Table 1: Comparison with the state-of-the-art pain estimation methods on the Biovid Heat Pain Database. The highest score is indicated in bold.
Method Modality CV Scheme Accuracy (%)
Werner et al. [60] ICPR 2014 EDA, ECG, EMG + Video 5-FOLD 80.6
Werner et al. [58] IEEE TAC 2016 Video LOSO 72.4
Kachele et al. [12] IEEE IJSTSP 2016 EDA, ECG, EMG LOSO 82.73
Lopez et al. [33] ACIIW 2017 EDA, ECG 10-FOLD 82.75
Lopez et al. [34] EMBC 2018 EDA LOSO 74.21
Thiam et al. [51] Sensors 2019 EDA LOSO 84.57
Wang et al. [56] EMBC 2020 EDA, ECG, EMG LOSO 83.3
Pouromran et al. [45] PLoS ONE 2021 EDA LOSO 83.3
Thiam et al. [52] Frontiers 2021 EDA, ECG, EMG LOSO 84.25
Phan et al. [44] IEEE Access 2023 EDA, ECG, EMG LOSO 84.8
1D CNN EDA 5-FOLD 77.2
R3D model Video 5-FOLD 72.9
Feature concatenation EDA + Video 5-FOLD 83.5
Vanilla transformer fusion EDA + Video 5-FOLD 87.8
Joint Multimodal Transformer (ours) EDA + Video 5-FOLD 89.1
@@ 4.2 Implementation Details
The audio and visual backbones are frozen to train the A-V fusion network and only train the whole transformer-based fusion model. Each of these backbones outputs deep feature vectors of dimension 512. These features are concatenated (to obtain the joint feature representation, a feature vector of dimension 1024) and then fed to an FC layer to reduce dimensionality to 512. Each of these features is fed to the fusion model, as we can see in Figure 2. At the output of each of the six cross-attention modules, the feature vector of dimension 512 is outputted. These six feature vectors are then stacked to form a sequence, which is then fed to a transformer self-attention block. This block dynamically selects and weighs these feature vectors. The final attended features are fed to a FC layer for final prediction. We perform a grid search to find the optimal learning rate and batch size for our fusion network. Thus, we use a list of learning rates: [$8\times 10^{-4}$, $6\times 10^{-4}$, $3\times 10^{-4}$] and an SGD optimizer to optimize the fusion network. We use a batch size of 32, and the maximum number of epochs is 5, with early stopping for fusion model selection. We select the best model among the learning rates listed before.
@@
For the modality fusion, the two feature vectors and the joint representation are fed to the joint transformer block, as show in Figure 3. The FC layers are removed from both backbones that were added in the backbone training phase. 512-dimensional feature vectors from visual and physiological backbones are obtained and fed to the joint multimodal transformer module. The backbones are frozen, and the joint transformer block is optimized using the ADAM optimizer with a learning rate of $5\times 10^{-6}$, and the batch size is set to 128.

### v3

@@ Abstract
Multimodal emotion recognition (MMER) systems typically outperform unimodal systems by leveraging the inter- and intra-modal relationships between, e.g., visual, textual, physiological, and auditory modalities. This paper proposes an MMER method that relies on a joint multimodal transformer (JMT) for fusion with key-based cross-attention. This framework can exploit the complementary nature of diverse modalities to improve predictive accuracy. Separate backbones capture intra-modal spatiotemporal dependencies within each modality over video sequences. Subsequently, our JMT fusion architecture integrates the individual modality embeddings, allowing the model to effectively capture inter- and intra-modal relationships. Extensive experiments [^1] Code: https://github.com/PoloWlg/Joint-Multimodal-Transformer-6th-ABAW. on two challenging expression recognition tasks – (1) dimensional emotion recognition on the Affwild2 dataset (with face and voice) and (2) pain estimation on the Biovid dataset (with face and biosensors) – indicate that our JMT fusion can provide a cost-effective solution for MMER. Empirical results show that MMER systems with our proposed fusion allow us to outperform relevant baseline and state-of-the-art methods.
@@
Figure 1: (Top) An illustration of the vanilla multimodal transformer fusion architecture in the case of two input sources, A and B. (Bottom) Our proposed JMT fusion (in red) relies on joint multimodal representations.
@@ 1 Introduction
Human-computer interaction is applied in a wide range of real-world scenarios, e.g., health care, the Internet of Things (IoT), and autonomous driving. Researchers have classified human emotions in different ways, most notably according to discrete categories, ordinal intensity levels, and the valence/arousal circumplex [1]. With the recent advancements in deep learning and sensor technologies, research in affective computing has evolved from lab-controlled to real-world (in the wild) scenarios. In the latter, human emotions are usually expressed over a broader spectrum beyond the six basic categorical expressions - anger, disgust, fear, happy, sad, and surprise [9]. Therefore, there is much interest in analyzing and modeling complex and subtle expressions of emotions in real-world scenarios. For instance, the spectrum of emotions can be formulated as dimensional emotional recognition (ER), where complex human emotions are represented along arousal (intensity) and valence (positiveness) axes.
@@
Several methods have been proposed for video-based ER and recurrent networks have been employed to capture the intra-modal temporal dependencies from video sequences [50, 28, 5]. Recently, attention-based methods have been introduced to extract features that are the most relevant to downstream tasks. Cross-attention-based methods have also been [48] employed to capture the inter-modal association between the audio, visual, and other modalities. Lu et al. [35] proposed ViLBERT, the seminal work in multimodal co-attention. Since then, many transformer-based cross-attention methods have been proposed [43, 57]. These methods, however, cannot effectively capture the intra-modal temporal dynamics. Further, they specialize in capturing the complementary information among the modalities but do not include a mechanism to explicitly capture the redundant information.
The proposed method introduces a third branch with the joint representation of the multiple modalities, as shown in Figure 1. By incorporating a joint representation branch, the model can access additional contextual information that may not be fully captured by cross-attention alone. Such a joint representation branch can help improve the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences, which helps mitigate the sensitivity of cross-attention to noisy inputs and improves the system’s overall performance.
@@
(1) This paper proposes a joint multimodal transformer (JMT) fusion architecture that leverages joint modality representations. It captures inter- and intra-modal information in videos using key-based cross-attention, and exploits the redundant and complementary associations among modalities. (2) An extensive set of experiments on two challenging emotion recognition datasets (pain estimation on BioVid and dimension valence-arousal assessment on Affwild2) indicate that our proposed JMT fusion architecture can outperform relevant baseline and state-of-the-art methods.
@@ 2.1 Multimodal Methods
MMER refers to integrating multiple sources of information (modalities) to improve the accuracy and robustness of automated emotion recognition systems at the expense of complexity. These modalities typically include visual, audio, textual, and physiological. The seminal work in multimodal deep learning was proposed by Ngiam et al. [39], where the features from the audio and visual modalities were extracted separately, and then autoencoders and Restricted Boltzmann Machines were used to feature fusion. Tzirakis et al. [55] proposed one of the early approaches for A-V fusion for dimensional emotion recognition, in which the visual features were extracted using a ResNet50 and the audio features were obtained using a 1D convolutional neural network (CNN). The modality-specific features were concatenated and fed to a recurrent net for simultaneous temporal modeling and modality fusion. An empirical study was presented by Juan et al. [42], where the authors studied the impact of fine-tuning multiple layers in a pretrained CNN for the visual modality.
A two-stream autoencoder with a long short-term memory (LSTM) network was proposed by Nguyen et al. [40] to jointly learn and compact representative features from the visual and audio modalities. A knowledge distillation-based approach was investigated by Schonevald et al. [50] for visual modality. For the audio modality, spectrograms were obtained and fed to a CNN model, and the two modalities were fused using a recurrent net. A novel self-distillation scheme was put forward by Deng et al. [6] to overcome the problem of noisy labels in a multitasking setting. A two-stream aural visual (TSAV) network was proposed by Kuhnke et al. [28], in which the audio features were extracted using a ResNet18, and the visual features were extracted using a 3D-CNN. The obtained embeddings were fed to a specially designed TSAV network for information fusion.
Pain estimation is one of the primary problems in affective computing. Researchers have proposed many multimodal datasets for the pain estimation task. The facial activity descriptors method for pain estimation was introduced by Werner et al. [59]. Dragomir et al. [7] propose a subject-independent method from facial images with a residual learning technique. A Sparse LSTM-based method was proposed by Zhi et al. [65] to solve the problem of vanishing gradients in temporal learning. Morabit et al. [38] proposed a data-efficient image transformer. To process multiscale electrodermal activity signals, a SE-Net-based network was proposed by Lu et al. [36]. Multimodal solutions to fuse the physiological and visual modalities were proposed by Werner et al. [59], Kachele et al. [13], and Zhu et al. [65]. Physiological signals are more discriminative for pain classification than the visual modality.
@@ 2.2 Attention-Based and Transformer Methods
Since its inception, attention models have shown extraordinary performance in many applications. These models have been extensively investigated for capturing the inter and intra-modal associations between the audio and visual modalities for tasks like action localization [32], A-V event localization [8], and multimodal emotion recognition [43]. An attention-based fusion mechanism was proposed by Zhang et al. [64], 3D-CNNs and 2D-CNNs were used to extract multi-features in the visual modality, and for the audio modality, a 2D-CNN was used to learn representation from spectrograms. Specialized scoring functions were used to re-weight the audio and visual features.
Recently, cross-modal attention has shown promising results because of its ability to model inter-modal relationships. Srinivas et al. [43] explored a transformer network with encoder layers, where cross-modal attention is used to fuse audio and visual features for continuous arousal/valence prediction in the wild. Tzirakis et al. [54] explored the idea of cross-attention in conjunction with self-attention. The authors proposed a transformer-based fusion architecture. Although the methods mentioned above have used cross-modal attention with transformers, they do not have any explicit mechanism to capture semantic relevance between the A-V features, particularly the intra-modal correlations. Zhang et al. [63] proposed a method for A-V fusion using leader-follower attentive fusion for continuous arousal/valence prediction. Attention weights are combined with the encoded visual and audio features. Cross attention presented in Praveen et al. [48] has shown a substantial increase in performance by using cross-correlation across the individual features. In contrast, our proposed method uses key-based cross-attention in multimodal transformers and explores the idea of feeding the joint A-V feature vector. By feeding the joint A-V feature representation, the proposed method effectively captures the inter- and intra-modal relationships simultaneously by interacting across itself and the other modalities.
@@
Huang et al. [11] investigated the idea of multi-head attention in transformer-based fusion architecture, which was further combined with LSTM to capture the high-level representations. Tran et al. [53] proposed a cross-modal transformer architecture that consisted of a multimodal cross-modal attention block, where the Queries were generated from one modality and the key values were generated from the other modality. Le et al. [31] put forward an end-to-end transformer-based fusion mechanism for multilabel multimodal emotion classification; the model consisted of three parts: i) three backbone networks for visual, audio, and textual feature extractor, ii) a transformer network for information fusion, and iii) classification network. Zhou et al. [66] proposed a transformer-based fusion scheme along with the temporal convolutional network (TCN); the audio and visual features were extracted using pretrained backbones followed by a TCN, the output of TCN was concatenated and fed to a transformer encoder block. A multilayer perceptron (MLP) was then used for the final prediction.
@@
All the aforementioned transformer-based fusion architectures primarily focus on intermodality correlation. In contrast, in addition to modeling the intermodality relationships to capture the complementarity between modalities, the proposed method explicitly feeds the joint (combined) features to the multimodal transformer to introduce redundancy. By incorporating this third joint representation branch, the proposed model can access enhanced contextual information that cross-attention might only partially capture. Doing this improves the model’s understanding of complex relationships between the input sequences. Further, the proposed method becomes more robust to noise or irrelevant information present in individual sequences. This third joint representation allows the model to dynamically focus on this newly introduced information in sequences where both modalities are simultaneously noisy. This helps mitigate the sensitivity of cross-attention to noisy inputs and improves the system’s overall performance.
@@ 3 Proposed Approach
The proposed method is a hierarchical fusion mechanism, where the intra-modality features are combined using transformer-based self-attention, and cross-modality features are fused using transformer-based cross-attention. Further, we feed a third joint representation to the joint transformer module (JTM) to enhance robustness. The $K$ (key matrix), $V$ (value matrix), and $Q$ (query matrix) vectors are shared among the six transformer blocks. In the end, the output of these six blocks is again fed to a transformer self-attention block to weigh the most relevant representations dynamically. The final prediction is made using fully connected (FC) layers.
@@ 3.1 Modality Specific Feature Extraction
In the first step, modality-specific features are extracted using backbones. The proposed method allows combining multiple backbones for each modality to improve system robustness. The extracted feature vectors from each backbone are fed to a transformer self-attention block. The combined feature vector represents the particular modality. For example, to capture information about a person’s emotional state, we can use an R(2+1)D CNN pretrained on the Kinetics-400 dataset [29] to extract visual features. For the audio modality, we could extract features using a ResNet18 [10] CNN with a GRU [4].
@@ 3.2 Multi-transformer Attention Fusion
Each representation is then fed to a specific encoder. Our model is composed of three different encoders, one for each type of feature ${\bm{f}}_{\mathbf{B}}$, ${\bm{f}}_{\mathbf{A}}$, and $\bm{f_{J}}$. Each encoder consists of a multi-head-self-attention (Eq. 2) followed by a fully connected feed-forward network. Residual connection and layer normalization are performed around both of these layers. The key is used to associate a sequence to a key value, the value matrix holds information that is ultimately used to compute the output of the attention mechanism, and the query matrix represents a set of vectors used to query the key-value pairs. The ${K}$, ${V}$, ${Q}$ matrix are calculated this way: ${K={X}{W_{K}}}$, ${Q={X}{W_{Q}}}$, ${V={X}{W_{V}}}$. ${X}$ corresponds to one of the sources of information and ${W_{K}}$, ${W_{Q}}$, ${W_{V}}$ are the weights of the key, query, value matrices respectively. The output values of the self-attention layers are given by:
${\mbox{Attention}(Q,K,V)=\mbox{softmax}({K}{Q^{T}}/{d_{k}})V}$ (2)
@@ 4.2 Implementation Details
The audio and visual backbones are frozen to train the A-V fusion network and only train the whole transformer-based fusion model. Each of these backbones outputs deep feature vectors of dimension 512. These features are concatenated (to obtain the joint feature representation, a feature vector of dimension 1024) and then fed to an FC layer to reduce dimensionality to 512. Each of these features is fed to the fusion model, as we can see in Figure 2. At the output of each of the six cross-attention modules, the feature vector of dimension 512 is outputted. These six feature vectors are then stacked to form a sequence, which is then fed to a transformer self-attention block. This block dynamically selects and weighs these feature vectors. The final attended features are fed to an FC layer for final prediction. We perform a grid search to find the optimal learning rate and batch size for our fusion network. Thus, we use a list of learning rates: [$8\times 10^{-4}$, $6\times 10^{-4}$, $3\times 10^{-4}$] and an SGD optimizer to optimize the fusion network. We use a batch size of 32, and the maximum number of epochs is 5, with early stopping for fusion model selection. We select the best model among the learning rates listed before.
@@
For the modality fusion, the two feature vectors and the joint representation are fed to the joint transformer block, as shown in Figure 3. The FC layers are removed from both backbones that were added in the backbone training phase. 512-dimensional feature vectors from visual and physiological backbones are obtained and fed to the joint multimodal transformer module. The backbones are frozen, and the joint transformer block is optimized using the ADAM optimizer with a learning rate of $5\times 10^{-6}$, and the batch size is set to 128.